# Mean-field theory of 1+1D $\mathbb{Z}_{2}$ lattice gauge theory with matter

### v1

@@ E.2 Fits of the entanglement entropy
$S_{\rm CFT}(x)=S_{0}+\frac{c}{6}\log{\left[\left(\frac{2L^{\prime}}{\pi}\right)\sin{\left(\frac{\pi x}{L^{\prime}}\right)}\right]}.$ ) . (62)
@@
For $\lambda\neq 0$ regime we see that $\tilde{S}(x)$ profiles are indeed almost completely flat for $1515<x<L-$ 15 for the regime away from criticality. When the system approaches the transition we note that the entanglement entropy obtains a signature curvature. As explained in the main text the central charge is lower at $c=0.5$ [46]. We note that our fit results are slightly worse as we frequently overshoot the correct value, see Fig. 15(c) and (d). However the qualitative picture is correct.
@@
We furthermore note that we only fitted the entanglement entropy profiles $\tilde{S}(x)$, in the range $1515<x<L-$ 15 in order to correctly capture the flat profiles in the parameter regimes away from criticality. At this point we would once again like to stress the fact that the CFT formula is only applicable exactly at the quantum criticality meaning that we expect poor fit results for generic parameters away from criticality [41].

### v2

@@ E.2 Fits of the entanglement entropy
[…]
@@
For $\lambda\neq 0$ regime we see that $\tilde{S}(x)$ profiles are indeed almost completely flat for 15<x<L1515<x<L-15 for the regime away from criticality. When the system approaches the transition we note that the entanglement entropy obtains a signature curvature. As explained in the main text the central charge is lower at $c=0.5$ [46]. We note that our fit results are slightly worse as we frequently overshoot the correct value, see Fig. 15(c) and (d). However the qualitative picture is correct.
@@
We furthermore note that we only fitted the entanglement entropy profiles $\tilde{S}(x)$, in the range 15<x<L1515<x<L-15 in order to correctly capture the flat profiles in the parameter regimes away from criticality. At this point we would once again like to stress the fact that the CFT formula is only applicable exactly at the quantum criticality meaning that we expect poor fit results for generic parameters away from criticality [41].